40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Sat Feb 20 16:20:33 2021, gautam, Update, General, Housekeeping + PRMI char IPPOSdeat.pdfTTdrift.pdfPOXnPOY.pdfPRMI.pdf
    Reply  Sun Feb 21 20:51:21 2021, rana, Update, General, Housekeeping + PRMI char 
       Reply  Tue Feb 23 00:10:05 2021, gautam, Update, General, Demod char part 1 noPreamp.pdfwithPreamp.pdf
          Reply  Wed Feb 24 11:53:24 2021, gautam, Update, General, Demod char part 2 noises_noPreamp.pdfnoises_withPreamp.pdf
             Reply  Wed Feb 24 12:11:08 2021, gautam, Update, General, Demod char part 3 noPreamp.pdfwithPreamp.pdfnoPreamp_whitened.pdfwithPreamp_whitened.pdf
                Reply  Thu Feb 25 16:50:53 2021, gautam, Update, General, PRMI sensing matrix PRMI1f_noArmssensMat.pdf
                   Reply  Thu Feb 25 20:37:49 2021, gautam, Update, General, Setting modulation frequency and checking IMC offset modFreq.pdfIMC_offset.pdf
Message ID: 15829     Entry time: Sat Feb 20 16:20:33 2021     Reply to this: 15831
Author: gautam 
Type: Update 
Category: General 
Subject: Housekeeping + PRMI char 

In prep to try some of these debugging steps, I did the following.

  1. ndscope updated from 0.7.9 to 0.11.3 on rossa. I've been testing/assisting the development for a few months now and am happy with it, and like the new features (e.g. PDF export). v0.7.9 is still available on the system so we can revert whenever we want.
  2. Arms locked on POX/POY, dither aligned to maximize TRX/TRY, normalization reset.
  3. PRMI locked, dither aligned to maximize POPDC.
  4. All vertex oplevs re-centered on their QPDs.

While working, I noticed that the annoying tip-tilt drift seems to be worse than it has been in the last few months. The IPPOS QPD is a good diagnostic to monitor stability of TT1/TT2. While trying to trend the data, I noticed that from ~31 Jan (Saturday night/Sunday morning local time), the IP-POS QPD segment data streams seem "frozen", see Attachment #1. This definitely predates the CDS crash on Feb 2. I confirmed that the beam was in fact incident on the IPPOS QPD, and at 1Y2/1Y3 that I was getting voltages going into the c1iscaux Acromag crate. All manner of soft reboots (eth1 network interface, modbusIOC service) didn't fix the problem, so I power cycled the acromag interface crate. This did the trick. I will take this opportunity to raise again the issue that we do not have a useful, reliable diagnsotic for the state of our Acromag systems. The problem seems to not have been with all the ADC cards inside the crate, as other slow ADC channels were reporting sensible numbers.

Anyways, now that the QPD is working again, you can see the drift in Attachment #2. I ran the dither alignment ~4 hours ago, and in the intervening time, the spot, which was previously centered on the AS camera CRT display, has almost drifted completely off (my rough calibration is that the spot has moved 5mm on the AS CCD camera). I was thinking we could try installing the two HAM-A coil drivers to control the TTs, this would allow us to rule out flaky electronics as the culprit, but I realize some custom cabling would be required, so maybe not worth the effort. The phenomenology of the drift make me suspect the electronics - hard for me to imagine that a mechanical creep would stop creeping after 3-4 hours? How would we explain the start of such a mechanical drift? On the other hand, the fact that the drift is almost solely in pitch lends support to the cause being mechanical. This would really hamper the locking efforts, the drift is on short enough timescales that I'd need to repeatedly go back and run the dither alignment between lock attempts - not the end of the world but costs ~5mins per lock attempt.


On to the actual tests: before testing the hardware, I locked the PRMI (no ETMs). In this configuration, I'm surprised to see that there is nearly perfect coherence between the MICH and PRCL error signals between 100Hz-1kHz 🤔 . When the AS55 demodulated signals are whitened prior to digitization (and then de-whitened digitally), the coherence structure changes. The electronics noise (measured with the PSL shutter closed) itself is uncorrelated (as it should be), and below the level of the two aforementioned spectra, so it is some actual signal I'm measuring there with the PRMI locked, and the coherence is on the light fields on the photodiode. So it would seem that I am just injecting a ton of AS55 sensing noise into the PRCL loop via the MICH->PRM LSC output matrix element. Weird. The light level on the AS55 photodiode has increased by ~2x after the September 2020 vent when we removed all the unused output optics and copper OMC. Nevertheless, the level isn't anywhere close to being high enough to saturate the ADC (confirmed by time domain signals in ndscope).

To get some insight into whether the whole RF system is messed up, I first locked the arm cavities with POX and POY as the error signals. Attachment #3 shows the spectra and coherence betweeen these two DoFs (and the dark noise levels for comparison). This is the kind of coherence profile I would expect - at frequencies where the loop gain isn't so high as to squish the cavity length noise (relative to laser frequency fluctuations), the coherence is high. Below 10 Hz, the coherence is lower than between 10-100 Hz because the OLG is high, and presumably, we are close to the sensing noise level. And above ~100 Hz, POX and POY photodiodes aren't sensing any actual relative frequency fluctuations between the arm length and laser frequency, so it's all just electronics noise, which should be incoherent.

The analogous plot for the PRMI lock is shown in Attachment #4. I guess this is telling me that the MICH sensing noise is getting injected into the PRCL error point between 100Hz-1kHz, where the REFL11 photodiode (=PRCL sensor) isn't dark noise limited, and so there is high coherence? I tuned the MICH-->PRM LSC output matrix element to minimize the height of a single frequency line driving the BS+PRM combo at ~313Hz in the PRCL error point. 

All the spectra are in-loop, the loop gain has not been undone to refer this to free-running noise. The OLGs themselves looked fine to me from the usual DTT swept sine measurements, with ~100 Hz UGF.

Attachment 1: IPPOSdeat.pdf  34 kB  | Hide | Hide all
IPPOSdeat.pdf
Attachment 2: TTdrift.pdf  23 kB  | Hide | Hide all
TTdrift.pdf
Attachment 3: POXnPOY.pdf  166 kB  Uploaded Sat Feb 20 21:43:42 2021  | Hide | Hide all
POXnPOY.pdf
Attachment 4: PRMI.pdf  226 kB  Uploaded Sat Feb 20 21:46:07 2021  | Hide | Hide all
PRMI.pdf
ELOG V3.1.3-